OpenStudy (surana):

What is the third term of (x+y)^9

OpenStudy (surana):

\[\left( x+y \right)^9\]

zepdrix (zepdrix):

\[\large\rm (x+y)^9\quad=\sum_{k=0}^9 \left(\begin{matrix}9 \\ \rm k\end{matrix}\right)x^{9-k}y^k\]Binomial Theorem, ya? Hmm no coefficients on the x or y, that makes this one a bit easier :))

zepdrix (zepdrix):

Oh you're still typing, hehe. I'll wait ^^

OpenStudy (surana):

Possible answers are: A: \[84x^6y\] B: \[84^7y^2\] C: \[36x^6y^3 \] D: \[36x^7y^2\]

OpenStudy (surana):

I apologize for taking a while, math's not my strong suit. Working on improving it though.

zepdrix (zepdrix):

We start counting from k=0, right? So we count 0, 1, 2... k=2 is going to be our third term. \[\large\rm third~term\quad=\left(\begin{matrix}9 \\ 2\end{matrix}\right)x^{9-2}y^2\]

OpenStudy (surana):

So I multiply that?

zepdrix (zepdrix):

Yes, simply the exponents. To deal with the 9 choose 2, you can rewrite it in terms of factorials, or if you don't mind the hassle, you can go down to the 10th row of pascals triangle

zepdrix (zepdrix):

This is one option :)\[\large\rm \left(\begin{matrix}n \\ \rm k\end{matrix}\right)=\frac{n!}{k!(n-k)!}\]

OpenStudy (surana):

So I put x and Y in the k and n sections and then multiply it by the 9 which is in the n! and k! sections?

zepdrix (zepdrix):

ummmm 0_o

OpenStudy (surana):

Oh, I messed up slightly, A is supposed to be \[84x^6y^3 \]

OpenStudy (surana):

And now I'm confused at where to put the numbers and variables in.

zepdrix (zepdrix):

Hmm ok, if Binomial Theorem is confusing, maybe you can do it another way. Just realize that the first thing in the brackets (x in this problem) starts at the largest power of 9, and counts down in each term. While the other thing starts at nothing, and counts up to 9. \[\rm (x+y)^9=\text{__}x^9y^0+\text{__}x^8y^1+\text{__}x^7y^2+\text{__}x^6y^3+...\]With a bunch more terms to follow, but we don't need those.

zepdrix (zepdrix):

I left a blank in front of each term, that is the coefficient coming from Pascal's Triangle.

OpenStudy (surana):

So it would be either A or C then.

zepdrix (zepdrix):

`third` means... three. yes? So are you sure you counted the `third` term correctly?

OpenStudy (surana):

So it'd be the one of the answers with 7 and 2 in it?

zepdrix (zepdrix):

yes.

OpenStudy (surana):

So with the big numbers, I multiply that one?

zepdrix (zepdrix):

What big numbers? I'm not sure what you're talking about. If you look at the 10th row of Pascal's Triangle, you'll see that the numbers go like this 1 9 36 84 126 126 84 36 9 1 These are the coefficients for our expansion ^

OpenStudy (surana):

Oh. Okay. So it'd be 36, then.

OpenStudy (surana):

So it'd be D!

zepdrix (zepdrix):

yes, good job. sorry the site keeps crashing on me.

OpenStudy (surana):

Thank you. I'll close this one now.